# Learning Primitive Embodied World Models: Towards Scalable Robotic Learning

### v2

@@ A.3 Key Lemma: Density via Topological Transitivity
By Axiom 2, $X$ is a compact metric space. By Axiom 1, $\mathcal{P}$ is finite, so $\langle\mathcal{P}\rangle_{+}$ is a finitely generated semigroup acting on $X$.
Suppose the action is topologically transitive: for any non-empty open sets $U,V\subset X$, there exists $f\in\langle\mathcal{P}\rangle_{+}$ such that $f(U)\cap V\neq\emptyset$. This is empirically valid: agents can reach any region from any other (up to environmental constraints).
Then, by the Birkhoff Transitivity Theorem for semigroup actions, the orbit $\langle\mathcal{P}\rangle_{+}\cdot x_{0}$ is dense in $X$ for some $x_{0}$.
Now consider path space $\mathcal{M}(X)$. Since $X$ is compact and connected, and $\mathcal{C}(\mathcal{P})$ generates piecewise-continuous paths with dense pointwise coverage, we can approximate any continuous path $\xi^{*}\in\mathcal{M}(X)$ by a sequence of composite primitives that “trace” $\xi^{*}$ within $\epsilon$-tubes.
More formally: for any $\epsilon>0$, partition $[0,T]$ into intervals $t_{i},t_{i+1}]$, and choose $f_{k_{i}}\in\mathcal{P}$ such that:
$d\left(f_{k_{i}}(x_{i}),\xi^{*}(t_{i+1})\right)<\epsilon.$
By continuity (Axiom 2) and compactness, such $f_{k_{i}}$ exist if $\mathcal{P}$ includes local reachability.
Thus, $\mathcal{C}(\mathcal{P})$ is dense in $\mathcal{M}(X)$. ∎
A.4 Cardinality Disparity via Computability Theory
Lemma A.2 (Cardinality Gap).
Under Axioms 13, we have:
$|\mathcal{P}|=n<\infty,\quad\text{but}\quad|\mathcal{D}_{\text{emb}}|=2^{\aleph_{0}}.$
Yet $\mathcal{C}(\mathcal{P})$ is dense in $\mathcal{D}_{\text{emb}}$.
Proof.
$|\mathcal{P}|=n$ by Axiom 1.
$\mathcal{D}_{\text{emb}}\subseteq C([0,T],X)$, the space of continuous functions from $[0,T]$ to $X$. Since $X\subset\mathbb{R}^{m}$, this space has cardinality $2^{\aleph_{0}}$ (uncountable).
However, $\mathcal{C}(\mathcal{P})$ is a countable union of finite-dimensional images, hence countable, but still dense (Lemma A.1). Thus:
$|\mathcal{D}_{\text{emb}}|\gg|\mathcal{P}|,\quad\text{yet }\overline{\mathcal{C}(\mathcal{P})}=\mathcal{D}_{\text{emb}}.$
This is analogous to: $\mathbb{Q}$ is countable, $\mathbb{R}$ is uncountable, but $\overline{\mathbb{Q}}=\mathbb{R}$. ∎
A.5 Categorical Compositionality: The Final Step
Define a category $\mathbf{Beh}$:
@@ A.5 Categorical Compositionality: The Final Step
    Objects: States $x\in X$.
@@
    Morphisms: Continuous paths $\xi:x\to y$.
@@
    Composition: Path concatenation.
Let $\mathbf{Prim}$ be the subcategory generated by $\mathcal{P}$: morphisms are finite composites of primitive transitions.
Theorem A.3 (Compositional Completeness).
Under Axioms 13, the inclusion functor:
$\iota:\mathbf{Prim}\hookrightarrow\mathbf{Beh}$
is dense in the categorical sense: every morphism in $\mathbf{Beh}$ is a limit of diagrams in $\mathbf{Prim}$.
Proof.
By Lemma A.1, $\mathcal{C}(\mathcal{P})$ is dense in $\mathcal{M}(X)$. By Lemma A.2, while $\mathrm{Hom}_{\mathbf{Prim}}(x,y)$ is countable, its closure equals $\mathrm{Hom}_{\mathbf{Beh}}(x,y)$. In the category of topological spaces and continuous paths, this implies that $\mathbf{Prim}$ is a dense generator of $\mathbf{Beh}$.
Hence, any behavior in $\mathcal{D}_{\text{emb}}$ is a colimit of primitive compositions. ∎
A.6 Conclusion: The Inevitability of Compositionality
We have shown that:
  • From three minimal, empirically grounded axioms,
  • Using only standard results from topology, dynamics, and category theory,
  • We derive that finite primitives can approximate any continuous embodied behavior.
The disparity $|\mathcal{D}_{\text{emb}}|\gg|\mathcal{P}|$ is not a limitation–it is a consequence of continuity and finiteness, and composition is the only way to bridge it.
Thus, compositionality is not a design choice.
It is a mathematical necessity of embodied intelligence.
@@ B.1 The Gen6D Method
Specifically, given a generated video sequence $V=\{I_{1},I_{2},\ldots,I_{T}\}$ and a reference video $V_{r}=\{I_{r1},I_{r2},\ldots,I_{rT^{\prime}}\}$ of the robotic gripper, we estimate the gripper’s 6-DoF pose using a model-free, RGB-based pose estimator: $\mathbf{R},\mathbf{T}=PE(V,V_{r})$, where $\mathbf{R}$ and $\mathbf{T}$ denote the rotation and translation matrices, respectively. In our implementation, we use Gen6D (Liu et al., 2022b) for its simplicity and generalization capability.
@@
After generating the future rollout $x_{1:T}^{\text{img}}$ for a primitive $a_{k}$, we extract a 6-DoF end-effector trajectory $\tau_{k}=\{p_{1},\ldots,p_{T}\}$ using 6-DoF pose estimation Liu et al. (2022b) and geometric transformation. These trajectories are then mapped into Cartesian control space and executed by the robot.
@@ Appendix C Details for Hierachical Plug-and-Play Long-Horizon Compositional Generalization
We leverage LoRA-based adaptation to enable reasoning and spatial grounding, following a Chain-of-LoRA (Liu et al., 2025b)-style architecture. The structure is detailed in Figure 8.
@@ D.1 Dataset Details
Our dataset consists of 7,326 simulated and 11,465 real-world primitives collected using a 5-camera synchronized setup with Deoxys (Zhu et al., 2022). By segmenting long-horizon tasks via keyframe indices, we extracted on average 5.8 primitives per session, achieving up to $29\times$ collection efficiency. As shown in Figure 14, cameras were arranged to maximize coverage of the workspace.
For annotation, 10% of the data was manually labeled, then used to fine-tune Qwen-VL 2.5-7B for auto-labeling the remainder, followed by light manual correction. To enhance visual generation quality, we mixed a small portion of diverse simulated data into the training set, including examples generated from RLBench and LIBERO. This hybrid strategy improves model generalization and dynamic realism, as confirmed in Section 5.3.
@@ D.2 Annotation Interface
Our annotation interface is designed to maximize labeling efficiency and consistency for multi-view robotic manipulation data. As shown in Figure 7 (see supplementary material), the system displays five synchronized views of the same primitive action simultaneously—ensuring that annotators can observe the full 3D context of each motion.
@@ E.1 Base Model Configuration
Our VLM planner is built upon the Qwen2.5-VL-7B-Instruct model (Team, 2025), a large-scale vision-language model capable of understanding and reasoning over both visual and textual inputs. Qwen-VL integrates a vision encoder based on the Vision Transformer (ViT) architecture to extract image features, which are then mapped into the language model’s embedding space via a cross-modal projector. The underlying LLM, with approximately 7 billion parameters, enables rich semantic understanding and multi-step reasoning, making it well-suited for complex task planning in robotic manipulation.
The model supports high-resolution image input and is pre-trained on a diverse corpus of image-text pairs, followed by supervised fine-tuning on instruction-following datasets. This training paradigm equips Qwen-VL with strong generalization capabilities across domains, including object recognition, spatial reasoning, and action sequence prediction—key competencies required for grounding high-level instructions into executable robotic plans.
@@ F.1 Video Generation Model Details
Our base model is DynamiCrafter (Xing et al., 2024), which has a 1.4B denoising network.
DynamiCrafter is architected upon the foundation of the open-source Text-to-Video (T2V) model, VideoCrafter (Chen et al., 2023; 2024a), and further integrates principles from the Text-to-Image (T2I) model, Stable-Diffusion-v2.1 (Rombach et al., 2022a). This hybrid approach allows DynamiCrafter to leverage the advanced capabilities of both text-conditioned video generation and stable image diffusion, forming a robust framework for animating open-domain images.
@@ G.1 Examples of Training Video Clips
We compare our framework against OpenVLA (Kim et al., 2024a), a state-of-the-art end-to-end vision-language-action model that maps raw instructions and observations directly to actions. We evaluate OpenVLA under two settings: (1) Zero-shot: The pre-trained OpenVLA model is deployed directly on our benchmark tasks without any task-specific finetuning. This setting evaluates its raw generalization capability across unseen objects and scenes. (2) Finetune: The model is finetuned on 100 task-specific demonstrations per task using supervised behavior cloning, following the same protocol as in our system’s policy training. This setting represents an upper-bound of OpenVLA’s performance under favorable conditions. We note that OpenVLA performs poorly in the zero-shot setting and fails completely on real-robot deployment, whereas our method maintains strong performance even without task-specific finetuning or retraining of the video model.
@@ H.2 Physically Accurate Generation
To provide a more comprehensive evaluation of physical fidelity in generated action sequences, we extend our metrics beyond standard video quality indicators (e.g., SSIM, PSNR, LPIPS) to include task- and embodiment-aware measures. In particular, we adopt the evaluation philosophy of PhysicsIQ (Motamed et al., 2025) and introduce the first Embodied Physical Consistency Score (EPiCS) tailored for robotic manipulation tasks. EPiCS is a fine-grained, human-in-the-loop metric that assesses whether generated motions respect fundamental physical and kinematic constraints in embodied environments. See a detailed description of EPiCS in Appendix J.
@@ H.3 Supplementary Comparisons
To further contextualize our method’s performance within the broader landscape of vision-based robotic planning, we provide supplementary comparisons on the RLBench benchmark (James et al., 2020b) against representative baselines from different paradigms. As shown in Table 7, our approach achieves significantly higher success rates on both Close Box and Open Drawer tasks compared to the VLM-based method VoxPoser (Huang et al., 2023) and the action-tokenization approach PerAct (Shridhar et al., 2022), where the performance of VoxPoser and PerAct are both extracted from the Colosseum paper (Pumacay et al., 2024). This performance gap highlights the advantage of our video diffusion-based policy generation, which produces temporally coherent and visually grounded action sequences, in contrast to methods that rely on discrete action token prediction or direct VLM-to-action mapping without explicit visual simulation. Notably, our method operates in a zero-shot manner on these tasks—without task-specific fine-tuning of the video generation model—further underscoring its strong generalization capability. These results complement the comprehensive evaluation in Table 1 and reinforce the effectiveness of our framework in handling common, yet non-trivial, manipulation primitives in simulation.
@@ H.4 Efficiency Comparison
The combination of low memory usage, high frame rate, and single-GPU compatibility makes our approach uniquely suitable for real-world embodied agents, where latency, cost, and hardware constraints are critical. It bridges the gap between high-fidelity simulation and deployable robot control—a key step toward scalable, real-time vision-to-action systems.
@@ H.5 Detailed Latency Analysis
It is important to emphasize that these measurements reflect unoptimized inference performance: no acceleration techniques such as reduced denoising steps, model distillation, tensor compilation, or hardware-specific optimization (e.g., GPU batching) have been applied. As such, the reported latencies represent a conservative lower bound on execution speed—or, equivalently, an upper bound on latency—under current implementation.
@@
H.6 Performance on Rotation-Intensive Tasks — An Illustration of Scalability
@@
As shown in Table 11, success rates on rotation-heavy tasks improve significantly—surpassing SOTA levels—while performance on non-rotational tasks remains stable, with no significant degradation.
H.7 Primitive-level Compositional Generalization
A core strength of our primitive-centric framework is its ability to achieve compositional generalization—the capacity to succeed on novel task configurations that were never explicitly demonstrated during training. This is realized through the modular design of primitives, which decouples the learning of actions (predicates) from the learning of object affordances, enabling the model to recombine known skills with new objects or known objects with new skills in a zero-shot manner.
We quantitatively evaluate this capability in Table 10, which reports success rates for executing unseen (predicate, object) combinations on a real robot. The results demonstrate strong generalization: for instance, the "pick" primitive, trained on objects like cups and boxes, successfully generalizes to the novel combination "pick jar" with an 80% success rate. Similarly, the "push" action generalizes effectively to cups and jars, despite these pairings being absent from the training data. Even for the more complex "open" action, the model transfers successfully to the unseen "open cup" task (70% success), showcasing its ability to adapt a skill to an object with significantly different geometry and affordances.
This generalization is enabled by two key factors. First, our video diffusion model, trained on densely annotated primitive data, learns rich, disentangled representations of action dynamics and object properties. This allows it to condition the generation of a "pick" motion on the visual features of a previously unseen jar. Second, the VLM-based primitive planner provides semantic and spatial grounding, correctly identifying the target object and generating a feasible subgoal configuration (e.g., "grasp the jar’s handle") even for unseen objects. The high success rates on these zero-shot combinations underscore that our approach moves beyond simple imitation learning, instead capturing the underlying compositional structure of embodied manipulation, a crucial step toward truly flexible and scalable robotic intelligence.
Predicate \ Object cup box drawer jar
pick 9/10 8/10 8/10
open 9/10 9/10 7/10
push 7/10 8/10 8/10 6/10
@@ H.7 Primitive-level Compositional Generalization
These results reinforce that our framework not only achieves competitive performance out-of-the-box but also enables practical, data-efficient improvement—making it well-suited for lifelong learning in dynamic environments.
@@ I.4 Long-Horizon Tasks
For each primitive in the sequence, the PEWM generates a short visual rollout conditioned on the current observation and the Start-Goal heatmap Guidance (SGG) priors. The 6-DoF end-effector trajectory is then extracted from this video using Gen6D and executed on the real robot. After each primitive’s execution, the robot’s state is updated, and the latest observation is fed back into the VLM planner. This creates an autoregressive loop where the planner can dynamically adjust the subsequent primitive sequence based on the outcome of the previous step—correcting for errors, handling unexpected disturbances (e.g., a moved object), or reacting to partial successes.

### v3

@@ A.2 Constructive Approximation Theorem
Since $X$ is compact and $\xi^{*}$ is continuous on $[0,T]$, it is uniformly continuous. Thus, for any $\epsilon>0$, there exists $\delta>0$ such that $|t-s|<\delta\Rightarrow\|\xi^{*}(t)-\xi^{*}(s)\|<\epsilon/2$.
Partition $[0,T]$ into $N$ intervals $t_{i},t_{i+1}]$ with $t_{i+1}-t_{i}<\min(\delta,T_{\max})$. For each segment $\xi^{*}_{i}=\xi^{*}|_{[t_{i},t_{i+1}]}$, define a local intent $u_{i}$ that describes the transition from $\xi^{*}(t_{i})$ to $\xi^{*}(t_{i+1})$ (e.g., “move end-effector from A to B”).
By Assumption A.1, $u_{i}$ maps to some $\tau_{k_{i}}\in\mathcal{T}$. By Assumption A.2, there exists a primitive instance $p_{i}$ (grounded via SGG or VLM) such that $p_{i}(0)=\xi^{*}(t_{i})$ and $\|p_{i}(t_{i+1}-t_{i})-\xi^{*}(t_{i+1})\|<\epsilon/2$. By continuity of $p_{i}$ and $\xi^{*}_{i}$, and since both are defined on an interval of length $\delta$, we have:
$\sup_{t\in[t_{i},t_{i+1}]}\|p_{i}(t-t_{i})-\xi^{*}(t)\|<\epsilon.$
Concatenating $\{p_{i}\}_{i=1}^{N}$ yields $\xi\in\mathcal{C}(\mathcal{T})$ with $d_{\infty}(\xi,\xi^{*})<\epsilon$. ∎
A.3 Cardinality Disparity and Data Efficiency
Let $|\mathcal{T}|=K\ll\infty$. The number of possible composed trajectories of length $N$ is at most $K^{N}$, which is countable. However, $\mathcal{M}_{T}(X)$ has cardinality $2^{\aleph_{0}}$ (uncountable). Despite this, Theorem A.4 shows that a countable set $\mathcal{C}(\mathcal{T})$ is dense in an uncountable space—exactly analogous to $\mathbb{Q}$ being dense in $\mathbb{R}$.
This implies:
@@ A.3 Cardinality Disparity and Data Efficiency
    Data efficiency: Learning $K$ primitives yields coverage of a dense subset of behaviors.
@@
    Generalization: Novel tasks are approximated by novel compositions, not novel primitives.
@@
    Scalability: Adding new objects/scenes requires no new primitive templates—only new grounding.
Thus, primitive-centric learning is not just practical—it is information-theoretically optimal for embodied intelligence under semantic constraints.
@@ B.1 The Gen6D Method
Specifically, given a generated video sequence $V=\{I_{1},I_{2},\ldots,I_{T}\}$ and a reference video $V_{r}=\{I_{r1},I_{r2},\ldots,I_{rT^{\prime}}\}$ of the robotic gripper, we estimate the gripper’s 6-DoF pose using a model-free, RGB-based pose estimator: $\mathbf{R},\mathbf{T}=PE(V,V_{r})$, where $\mathbf{R}$ and $\mathbf{T}$ denote the rotation and translation matrices, respectively. In our implementation, we use Gen6D (gen6d) for its simplicity and generalization capability.
@@
After generating the future rollout $x_{1:T}^{\text{img}}$ for a primitive $a_{k}$, we extract a 6-DoF end-effector trajectory $\tau_{k}=\{p_{1},\ldots,p_{T}\}$ using 6-DoF pose estimation gen6d and geometric transformation. These trajectories are then mapped into Cartesian control space and executed by the robot.
@@ Appendix C Details for Hierachical Plug-and-Play Long-Horizon Compositional Generalization
We leverage LoRA-based adaptation to enable reasoning and spatial grounding, following a Chain-of-LoRA (liu2025videomind)-style architecture. The structure is detailed in Figure 8.
@@ D.1 Dataset Details
Our dataset consists of 7,326 simulated and 11,465 real-world primitives collected using a 5-camera synchronized setup with Deoxys (zhu2022viola). By segmenting long-horizon tasks via keyframe indices, we extracted on average 5.8 primitives per session, achieving up to $29\times$ collection efficiency. As shown in Figure 14, cameras were arranged to maximize coverage of the workspace.
For annotation, 10% of the data was manually labeled, then used to fine-tune Qwen-VL 2.5-7B for auto-labeling the remainder, followed by light manual correction. To enhance visual generation quality, we mixed a small portion of diverse simulated data into the training set, including examples generated from RLBench and LIBERO. This hybrid strategy improves model generalization and dynamic realism, as confirmed in Section 3.
@@ D.2 Annotation Interface
Our annotation interface is designed to maximize labeling efficiency and consistency for multi-view robotic manipulation data. As shown in Figure 7 (see supplementary material), the system displays five synchronized views of the same primitive action simultaneously–ensuring that annotators can observe the full 3D context of each motion.
@@ E.1 Base Model Configuration
Our VLM planner is built upon the Qwen2.5-VL-7B-Instruct model (qwen2.5-VL), a large-scale vision-language model capable of understanding and reasoning over both visual and textual inputs. Qwen-VL integrates a vision encoder based on the Vision Transformer (ViT) architecture to extract image features, which are then mapped into the language model’s embedding space via a cross-modal projector. The underlying LLM, with approximately 7 billion parameters, enables rich semantic understanding and multi-step reasoning, making it well-suited for complex task planning in robotic manipulation.
The model supports high-resolution image input and is pre-trained on a diverse corpus of image-text pairs, followed by supervised fine-tuning on instruction-following datasets. This training paradigm equips Qwen-VL with strong generalization capabilities across domains, including object recognition, spatial reasoning, and action sequence prediction–key competencies required for grounding high-level instructions into executable robotic plans.
@@ F.1 Video Generation Model Details
Our base model is DynamiCrafter (xing2024dynamicrafter), which has a 1.4B denoising network.
DynamiCrafter is architected upon the foundation of the open-source Text-to-Video (T2V) model, VideoCrafter (chen2023videocrafter1; chen2024videocrafter2), and further integrates principles from the Text-to-Image (T2I) model, Stable-Diffusion-v2.1 (Rombach_2022_CVPR). This hybrid approach allows DynamiCrafter to leverage the advanced capabilities of both text-conditioned video generation and stable image diffusion, forming a robust framework for animating open-domain images.
@@ G.1 Examples of Training Video Clips
We compare our framework against OpenVLA (kim2024openvla), a state-of-the-art end-to-end vision-language-action model that maps raw instructions and observations directly to actions. We evaluate OpenVLA under two settings: (1) Zero-shot: The pre-trained OpenVLA model is deployed directly on our benchmark tasks without any task-specific finetuning. This setting evaluates its raw generalization capability across unseen objects and scenes. (2) Finetune: The model is finetuned on 100 task-specific demonstrations per task using supervised behavior cloning, following the same protocol as in our system’s policy training. This setting represents an upper-bound of OpenVLA’s performance under favorable conditions. We note that OpenVLA performs poorly in the zero-shot setting and fails completely on real-robot deployment, whereas our method maintains strong performance even without task-specific finetuning or retraining of the video model.
@@ H.2 Physically Accurate Generation
To provide a more comprehensive evaluation of physical fidelity in generated action sequences, we extend our metrics beyond standard video quality indicators (e.g., SSIM, PSNR, LPIPS) to include task- and embodiment-aware measures. In particular, we adopt the evaluation philosophy of PhysicsIQ (motamed2025generative) and introduce the first Embodied Physical Consistency Score (EPiCS) tailored for robotic manipulation tasks. EPiCS is a fine-grained, human-in-the-loop metric that assesses whether generated motions respect fundamental physical and kinematic constraints in embodied environments. See a detailed description of EPiCS in Appendix J.
@@ H.3 Supplementary Comparisons
To further contextualize our method’s performance within the broader landscape of vision-based robotic planning, we provide supplementary comparisons on the RLBench benchmark (james2019rlbench) against representative baselines from different paradigms. As shown in Table 8, our approach achieves significantly higher success rates on both Close Box and Open Drawer tasks compared to the VLM-based method VoxPoser (huang2023voxposer) and the action-tokenization approach PerAct (shridhar2022peract), where the performance of VoxPoser and PerAct are both extracted from the Colosseum paper (pumacay2024colosseum). This performance gap highlights the advantage of our video diffusion-based policy generation, which produces temporally coherent and visually grounded action sequences, in contrast to methods that rely on discrete action token prediction or direct VLM-to-action mapping without explicit visual simulation. Notably, our method operates in a zero-shot manner on these tasks–without task-specific fine-tuning of the video generation model–further underscoring its strong generalization capability. These results complement the comprehensive evaluation in Table 1 and reinforce the effectiveness of our framework in handling common, yet non-trivial, manipulation primitives in simulation.
@@ H.4 Efficiency Comparison
The combination of low memory usage, high frame rate, and single-GPU compatibility makes our approach uniquely suitable for real-world embodied agents, where latency, cost, and hardware constraints are critical. It bridges the gap between high-fidelity simulation and deployable robot control–a key step toward scalable, real-time vision-to-action systems.
@@ H.5 Detailed Latency Analysis
It is important to emphasize that these measurements reflect unoptimized inference performance: no acceleration techniques such as reduced denoising steps, model distillation, tensor compilation, or hardware-specific optimization (e.g., GPU batching) have been applied. As such, the reported latencies represent a conservative lower bound on execution speed–or, equivalently, an upper bound on latency–under current implementation.
@@
H.6 Performance on Rotation-Intensive Tasks – An Illustration of Scalability
@@
As shown in Table 11, success rates on rotation-heavy tasks improve significantly–surpassing SOTA levels–while performance on non-rotational tasks remains stable, with no significant degradation.
@@
These results reinforce that our framework not only achieves competitive performance out-of-the-box but also enables practical, data-efficient improvement–making it well-suited for lifelong learning in dynamic environments.
@@ I.4 Long-Horizon Tasks
For each primitive in the sequence, the PEWM generates a short visual rollout conditioned on the current observation and the Start-Goal heatmap Guidance (SGG) priors. The 6-DoF end-effector trajectory is then extracted from this video using Gen6D and executed on the real robot. After each primitive’s execution, the robot’s state is updated, and the latest observation is fed back into the VLM planner. This creates an autoregressive loop where the planner can dynamically adjust the subsequent primitive sequence based on the outcome of the previous step–correcting for errors, handling unexpected disturbances (e.g., a moved object), or reacting to partial successes.
@@
Appendix L Other Promising Applications
The modular and explainable nature of our framework unlocks several other promising applications.
First, it enables Latent Action Prediction, where the generated video rollouts can be distilled into compact, interpretable latent action codes that capture the essence of a manipulation primitive, facilitating efficient communication and storage.
Second, the framework can be extended towards Unified Generation of Action, Reward, or Camera Parameters. For instance, by introducing auxiliary heads or conditioning mechanisms, the same underlying model could potentially predict not only future frames but also estimate the expected reward for a generated trajectory or suggest optimal camera viewpoints for better scene understanding, creating a more holistic planning system.
Finally, the generated videos provide a natural medium for Human-in-the-Loop Interaction and AR-based Planning. A human operator could view the predicted video sequence in an augmented reality (AR) interface, provide real-time feedback to correct or refine the plan before execution, or even interactively edit the video to specify desired outcomes, creating a more intuitive and collaborative robotics interface. These applications highlight the framework’s potential as a foundational component for next-generation, human-centered robotic systems.
Appendix M The Use of Large Language Models (LLMs)
In the preparation of this paper, large language models (LLMs) were used as an assistive tool to aid in the polishing and refinement of written content. Specifically, LLMs were employed to: Improve clarity, grammar, and fluency of draft text; Suggest rephrasings for complex or awkwardly worded sentences; Assist in generating concise summaries or transitional phrases where needed.
The core research ideas, experimental design, data analysis, interpretation of results, and final structuring of arguments were conceived and executed solely by the human authors. LLMs did not contribute to the generation of novel research hypotheses, methodological decisions, or scientific conclusions. All factual claims, citations, and technical content were verified and validated by the authors.